Please help me with this question.
When you add three to a number and then divide the sum by seven, you get the same result as when you add three to the same number and then divide the sum by twelve. What is the number?
Thank you for your help.

Answer by TimothyLamb(4379)

---
(3 + x)/7 = (3 + x)/12
12(3 + x) = 7(3 + x)
36 + 12x = 21 + 7x
5x = -15
x = -3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php